OpenStudy (anonymous):

x^2+3x+2≤0 can you please solve this problem and show me how to do it thanks

OpenStudy (akshay_budhkar):

ok lets first solve the quadratic we have \[x^2+3x+2=x^2+2x+x+2=x(x+2)+1(x+2)=(x+1)(x+2)\]

OpenStudy (akshay_budhkar):

so \[(x+1)(x+2)\le0 \therefore -2\le x \le-1\]